OpenStudy (anonymous):

Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 3 liters of natural oil for every 4 liters of synthetic oil. If 192 liters of natural oil are used, how many liters of Petrolyn oil will be made?

OpenStudy (anonymous):

@HelennnnnnC @chetan552 @StudyGurl14

OpenStudy (studygurl14):

3 L of natural oil + 4 liters of synthetic = 7 L Petrolyn oil \[\frac{ 3 }{ 7 } = \frac{ 192 }{ P }\] \[3P = 7(192)\] \[3P = 1344\] \[P = 448\] 448 liters of Petrolyn oil

OpenStudy (studygurl14):

I'm positive 448 L is the answer, @chetan552

OpenStudy (anonymous):

Sorry @StudyGurl14 is correct...

OpenStudy (studygurl14):

It's fine @chetan552